# Reflecting on Empirical and Sustainability Aspects of Software Engineering Research in the Era of Large Language Models

### v1

@@ 5. Future Plans
In future, we plan to expand our analysis to other important empirical aspects, such as the clarity of reporting on prompts and datasets used. Further, we aim to tackle additional ethical and social factors, such as whether the cost and infrastructure demands of LLMs are contributing to a digital divide in SE research, particularly between well-resourced and low-resourced institutions.
6. Conclusions
Our study provides an initial reflection on LLM-based empirical SE studies with respect to crucial empirical and ethical aspects. We hope that our results encourage the community to take up the open challenges and onboard the suggestions highlighted in our work.
@@ Acknowledgements.
Max Hort is supported by the European Union through the Horizon Europe Marie Skłodowska-Curie Actions (#101151798).References
@@
- Liang et al. (2024) Jenny T. Liang, Carmen Badea, Christian Bird, Robert DeLine, Denae Ford, Nicole Forsgren, and Thomas Zimmermann. 2024. Can GPT-4 Replicate Empirical Software Engineering Research? Proc. ACM Softw. Eng. 1, FSE, Article 60 (2024), 24 pages. doi:10.1145/3660767
@@
- Nie et al. (2025a) Yuqing Nie, Chong Wang, Kailong Wang, Guoai Xu, Guosheng Xu, and Haoyu Wang. 2025a. Decoding Secret Memorization in Code LLMs Through Token-Level Characterization. In 2025 IEEE/ACM 47th International Conference on Software Engineering (ICSE). 2880–2892. doi:10.1109/ICSE55347.2025.00229
- Nie et al. (2025b) Yuqing Nie, Chong Wang, Kailong Wang, Guoai Xu, Guosheng Xu, and Haoyu Wang. 2025b. Decoding Secret Memorization in Code LLMs Through Token-Level Characterization. In 2025 IEEE/ACM 47th International Conference on Software Engineering (ICSE). IEEE, 2880–2892.
@@
- Sallou et al. (2024) June Sallou, Thomas Durieux, and Annibale Panichella. 2024. Breaking the Silence: the Threats of Using LLMs in Software Engineering. In Proceedings of the 2024 ACM/IEEE 44th International Conference on Software Engineering: New Ideas and Emerging Results (ICSE-NIER’24). Association for Computing Machinery, 102–106. doi:10.1145/3639476.3639764
@@
- Yang et al. (2024a) Zhou Yang, Zhipeng Zhao, Chenyu Wang, Jieke Shi, Dongsun Kim, Donggyun Han, and David Lo. 2024a. Unveiling Memorization in Code Models. In Proceedings of the IEEE/ACM 46th International Conference on Software Engineering (Lisbon, Portugal) (ICSE ’24). Association for Computing Machinery, New York, NY, USA, Article 72, 13 pages. doi:10.1145/3597503.3639074
- Yang et al. (2024b) Zhou Yang, Zhipeng Zhao, Chenyu Wang, Jieke Shi, Dongsun Kim, Donggyun Han, and David Lo. 2024b. Unveiling memorization in code models. In Proceedings of the IEEE/ACM 46th International Conference on Software Engineering. 1–13.

### v2

@@ 4. Conclusions & Future Plans
In future work, we plan to deepen our analysis to cover additional empirical aspects, such as the clarity and rigour of reporting on prompts and datasets used. Further, we aim to tackle additional ethical and social factors, including whether the costs and infrastructure demands of LLMs are enabling a digital divide in the field, particularly between well-resourced and low-resourced institutions.
Ultimately, we aim for our work to contribute to the emerging efforts by the wider community towards developing unified, actionable guidelines for the responsible and sustainable use of LLMs in software engineering research.
@@ Acknowledgements.
We thank the anonymous participants of our survey. Max Hort is supported by the European Union through the Horizon Europe Marie Skłodowska-Curie Actions (#101151798). Aldeida Aleti is supported by the Australian Research Council, DP210100041.References
@@
- Baltes et al. (2025) Sebastian Baltes, Florian Angermeir, Chetan Arora, Marvin Muñoz Barón, Chunyang Chen, Lukas Böhme, Fabio Calefato, Neil Ernst, Davide Falessi, Brian Fitzgerald, Davide Fucci, Marcos Kalinowski, Stefano Lambiase, Daniel Russo, Mircea Lungu, Lutz Prechelt, Paul Ralph, Rijnard van Tonder, Christoph Treude, and Stefan Wagner. 2025. Guidelines for Empirical Studies in Software Engineering involving Large Language Models. arXiv:2508.15503 [cs.SE] https://arxiv.org/abs/2508.15503
@@
- Nie et al. (2025) Yuqing Nie, Chong Wang, Kailong Wang, Guoai Xu, Guosheng Xu, and Haoyu Wang. 2025. Decoding Secret Memorization in Code LLMs Through Token-Level Characterization. In Proceedings of the IEEE/ACM 47th International Conference on Software Engineering (Ottawa, Ontario, Canada) (ICSE ’25). IEEE Press, 2880–2892. doi:10.1109/ICSE55347.2025.00229
@@
- Sainz et al. (2023) Oscar Sainz, Jon Campos, Iker García-Ferrero, Julen Etxaniz, Oier Lopez de Lacalle, and Eneko Agirre. 2023. NLP Evaluation in trouble: On the Need to Measure LLM Data Contamination for each Benchmark. In Findings of the Association for Computational Linguistics: EMNLP 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 10776–10787. doi:10.18653/v1/2023.findings-emnlp.722
- Sallou et al. (2024) June Sallou, Thomas Durieux, and Annibale Panichella. 2024. Breaking the Silence: the Threats of Using LLMs in Software Engineering. In Proceedings of the 2024 ACM/IEEE 44th International Conference on Software Engineering: New Ideas and Emerging Results (Lisbon, Portugal) (ICSE-NIER’24). Association for Computing Machinery, New York, NY, USA, 102–106. doi:10.1145/3639476.3639764
@@
- Yang et al. (2024) Zhou Yang, Zhipeng Zhao, Chenyu Wang, Jieke Shi, Dongsun Kim, Donggyun Han, and David Lo. 2024. Unveiling Memorization in Code Models. In Proceedings of the IEEE/ACM 46th International Conference on Software Engineering (Lisbon, Portugal) (ICSE ’24). Association for Computing Machinery, New York, NY, USA, Article 72, 13 pages. doi:10.1145/3597503.3639074